# Augmentation-based Unsupervised Cross-Domain Functional MRI Adaptation for Major Depressive Disorder Identification

### v1

@@ I Introduction
Figure 1: Overview of our augmentation-based unsupervised cross-domain fMRI adaptation (AUFA) framework, including (1) fMRI feature extraction for brain functional connectivity networks using a Transformer module, (2) domain adaptation module for feature alignment between source and target domains, (3) augmentation-based self-optimization module, and (4) a classification module. The target data are fed into the model that has been jointly optimized by $L_{C}$, $L_{M}$ and $L_{A}$ for direct prediction.
@@
Major depressive disorder (MDD) is one of the most common mental health disorders, leading to symptoms such as extremely low mood, self-denial, feelings of helplessness, and impaired cognitive function, thus significantly impacting the patient’s daily life and work [1]. Currently, the diagnosis of MDD mainly relies on the patient’s symptom assessment, physical examination, and psychological scales, which still involves subjectivity and risk of misdiagnosis [2, 3]. Therefore, accurate detection of MDD is crucial in helping clinicians to formulate timely intervention plans.
@@
Resting-state functional magnetic resonance imaging (rs-fMRI), as a non-invasive and non-radioactive technique, can assess changes in neural activity across the brain over time. It has been widely used in computer-aided analysis of brain disorders using learning-based methods [4, 5]. Currently, machine learning/deep learning techniques generally focus on using functional connectivity networks (FCN) derived from blood-oxygen-level dependent (BOLD) time series of rs-fMRI data to capture interactions between brain regions-of-interest (ROIs) [6]. Traditional machine learning methods often extract handcrafted features based on FCN for brain disease detection. However, these methods require prior knowledge of FCN [7], and shallow features usually result in poor performance in classification tasks.
@@
With the development of deep learning, numerous methods employ graph neural networks to automatically learn high-level feature representations of fMRI data for automatic diagnosis of MDD [8, 9]. For example, Li et al[10] extracted topological features from brain functional connectivity constructed based on fMRI data, and utilized a graph neural network (GNN) to identify MDD. Jun et al[11] combined valid connectivity features from rs-fMRI with non-imaging phenotypic features, and the combined features were further learned by a graph convolutional network (GCN). However, these methods only use a single site from the MDD dataset for experiments, which does not guarantee the generalization ability of the model. Multi-site rs-fMRI data provide more data for learning-based methods to train reliable models [12, 13]. However, fMRI acquired from different sites often exhibit significant heterogeneity due to equipment differences and variations in data acquisition conditions [14], which presents a challenge for diagnosing brain diseases.
@@
Domain adaptation helps reduce the data distribution difference between the source and target domains [15]. However, when the gap between source and target domains is large, coupled with the size limitation of fMRI datasets, it becomes easier for the model to overfit on the source domain  [16]. That is, the model is too fine-grained in capturing the source domain features and ignores the generalization of the universal laws. Intuitively, augmenting the target domain data can enhance the model’s generalization ability to the target domain and avoid over-reliance on specific source domain data [17, 18].
@@
To this end, we propose a new augmentation-based unsupervised cross-domain fMRI adaptation (AUFA) framework for automated diagnosis of MDD. As shown in Fig. 1, we first construct FCNs for source/target domain subjects based on fMRI time series. Then, the Transformer encoder is utilized to learn fMRI features with spatial attention. At the same time, we introduce the maximum mean difference constraint to align the fMRI features learned from the source and target domains, aiming for the Transformer encoder to learn the shared features between domains. In addition, to further improve the model’s generalization ability to the target domain, we optimize the model by augmenting the data in the target domain to prevent the model from overfitting on the source domain. AUFA can significantly reduce the impact of data distribution differences across different sites through the supervision of source domain labels and adaptation of data between domains. Experiments on rs-fMRI data from 1,089 subjects demonstrate the excellent performance of AUFA in MDD identification compared with competing methods.
The major contributions of this work are summarized below.
@@
  • A novel augmentation-based unsupervised cross-domain fMRI adaptation (AUFA) framework is proposed for automated diagnosis with resting-state fMRI data.
@@
  • The proposed AUFA introduces the maximum mean difference constraint to align fMRI features from different sites, which can effectively reduce the data distribution differences between source and target domains.
@@
  • AUFA can effectively capture the attention relationship between different brain regions to automatically learn fMRI features and obtain discriminative functional connectivity.
  • Cross-domain disease detection experiments at 5 fMRI imaging sites show that our method is superior to several state-of-the-art methods.
The remainder of this paper is organized as follows. Section II provides a brief review of related studies. Section III introduces the data acquisition and pre-processing. Section IV introduces the details of the proposed method. In Section V, we present the experimental setup, competing methods, and results. Section VI discusses key components of the proposed method and list several limitations of the current work. Finally, we summarize the paper in Section VII.
@@ II-A Learning-based Methods for Functional MRI Analysis
Deep learning, as a powerful machine learning method, has been widely used in fMRI analysis [19, 20]. Compared with traditional machine learning methods, deep learning models can better handle high-dimensional, nonlinear and complex fMRI data with better modeling capabilities and generalization performance [21]. Numerous studies have used multilayer perceptron or convolutional neural network models to extract features of FCNs constructed based on fMRI data [22]. Graph neural networks, as a class of deep learning algorithms dedicated to processing graph data, have also been widely used for the analysis of FCN based on fMRI, such as graph convolutional networks (GCNs) [23]. GCNs learn node representations by aggregating features of nodes and their neighbors through convolutional operations. Yao et al.et al[24] proposed a mutual multi-scale triplet GCN to analyze brain functional and structural connectivity for brain disease diagnosis. Noman et al.et al[25] designed a novel graph autoencoder (GAE) architecture based on GCN for MDD detection. However, these methods usually overlook the different contributions of different brain regions to the classification task. Transformer is an attention-based deep learning model originally proposed for natural language processing tasks such as machine translation [26]. It introduces a self-attention mechanism that enables modeling the dependencies between different positions in an input sequence. In recent years, the Transformer model has also been applied as a powerful sequence modeling tool for capturing attention relationships between brain regions in fMRI data analysis [27, 28]. In this work, we attempt to leverage the Transformer in unsupervised domain adaptation for multi-site fMRI feature extraction. Considering that the strong backbone of Transformer increases the risk of overfitting, we augment the target domain data and improve the model generalization over the target domain through a self-optimization module.
@@ II-B Domain Adaptation of Functional MRI Data
The multi-site fMRI data greatly increased the sample size for the model training. However, due to distributional differences among data from different sites, the training data may not accurately reflect the distribution of the test data, leading to a degradation in model performance [29, 14]. The domain adaptation approach is an important strategy for solving this problem. Currently, domain adaptation methods can be classified into three categories based on the availability of target domain labels [30]: (1) supervised domain adaptation, where a small amount of labeled data in the target domain can be used for model training [31]; (2) semi-supervised domain adaptation, where a small amount of labeled data and redundant unlabeled data in the target domain are available during the training process [32]; and (3) unsupervised domain adaptation, where only unlabeled target data can be used to train the adaptive model [33]. In this paper, we focus on unsupervised domain adaptation.
In recent years, various domain adaptation methods have been widely used in the field of computer vision [34]. For example, Zhou et al[35] proposed a new Adversarial Distribution Adaptation Network (ADAN) to jointly reduce the global and local distributional differences between different domains to learn domain invariant representations. Li et al[36] proposed a Progressive Adaptation of Subspace (PAS) method to gradually obtain reliable pseudo-labels for model training. Liu et al[37] developed a novel unsupervised domain adaptation framework that ensures the preservation of the source data distribution to guide target structure learning in a semi-supervised clustering manner. However, multi-site adaptation studies for fMRI time series are still limited. Some studies based on multi-site fMRI data have simply combined fMRI data from different sites for training [38], i.e., assuming that they have the same data distribution, ignoring data heterogeneity between sites. Even though some studies attempt to reduce the distributional differences of fMRI data through domain adaptation methods, they usually ignore the problem that the model is likely to be overfitted on the source domain. For example, Wang et al[39] proposed an unsupervised graph domain adaptation network for the diagnosis of autism spectrum disorders, which consists of three main components: graph isomorphism encoder, progressive feature alignment, and unsupervised information matrix regularization. Lee et al[29] designed a new domain generalization meta-learning model for the diagnosis of brain diseases based on rs-fMRI. The aim of this paper is to study unsupervised domain adaptation for multi-site fMRI data from two perspectives, i.e., reducing inter-domain distributional differences and mitigating the risk of model overfitting.
TABLE I: Phenotypic information of 1089 subjects from REST-meta-MDD consortium [40]. MDD: Major depressive disorder; NC: Normal control; M/F: Male/Female. The age information is reported as mean±$\pm$±standard deviation.
@@ III-A Data Acquisition
This work involved rs-fMRI data from the top five largest sites in the imaging center of the REST-meta-MDD consortium [40], including SITE 20, SITE 21, SITE 25, SITE 1, and SITE 9. There were a total of 1,089 participants across the five sites, including 581 MDD patients and 508 normal controls (NCs). The rs-fMRI data from SITE 20, SITE 21 and SITE 1 were acquired by Siemens Verio 3T with the following parameters: echo time (TE) $=30\,ms$, repetition time (TR) $=2,000\,ms$, flip angle $=90$. Particularly, at SITE 20, the thickness/gap was $3.0\,mm/1.0\,mm$ with $32$ slice; at SITE 21, the thickness/gap was $3.5\,mm/0.7\,mm$ with $32$ slice, and at SITE 1, the thickness/gap was $4.0\,mm/0.8\,mm$ with $32$ slice. The rs-fMRI data from SITE 25 were obtained by Siemens Verio 3T with the following parameters: TE $=25\,ms$, TR $=2,000\,ms$, flip angle$=90$, thickness/gap $=4.0\,mm/0\,mm$, slice number $=36$. The rs-fMRI data for SITE 9 were obtained by GE Discovery MR750 3.0T with the following parameters: TE $=25\,ms$, TR $=2,000\,ms$, flip angle$=90$, thickness/gap = $3.0\,mm/1.0\,mm$, slice number $=35$. More detailed demographic information is reported in Table I.
@@ III-B Image Pre-Processing
We pre-processed the initial rs-fMRI data through the Data Processing Assistant for Resting-State fMRI (DPARSF) [41] pipeline, including the following steps: (1) discard the first ten volumes, (2) head motion correction, (3) space normalization and smoothing, (4) interference signal removal, (5) bandpass filtering. Finally, we extracted an average rs-fMRI time series for 116 predefined regions of interest (ROI) based on Automated Anatomical Labeling (AAL) [42] atlas.
@@ IV-A Problem Formulation
In this work, we focus on using cross-domain rs-fMRI data for the MDD identification task, aiming to solve an interesting, realistic, but challenging unsupervised domain adaptation problem, i.e., utilizing information learned from the source domain to assist the classification of a completely unlabeled target domain. We first introduce some notation that will be used in this work. Let $X^{d}$ be the input (graphs constructed based on fMRI data) space and $Y^{d}$ be the target (label) space, where $d$ identifies either the source domain $s$ or the target domain $t$. The source domain is defined as $S$ = $\{X^{s},Y^{s}\}$, containing the feature set $X^{s}$ and the corresponding label set $Y^{s}$. The target domain is defined accordingly as $T$ = $\{X^{t},Y^{t}\}$ with feature set $X^{t}$ and label set $Y^{t}$. Even though the two domains have the same task (they share the same label space $Y^{s}$ = $Y^{t}$), their data distribution is somewhat different, i.e., $P^{s}(X)\neq P^{t}(X)$. In the framework of unsupervised domain adaptation, the challenge is how to learn a feature extractor $\phi:X\rightarrow Z$ that leverages the abundant labeling information available in the source domain. This extractor maps the initial input to a feature space $Z$, aiming to minimize the data distribution between the source and target domains, i.e., $P^{s}(\phi(X))\approx P^{t}(\phi(X))$.
@@ IV-B Method
We propose an augmentation-based unsupervised cross-domain fMRI adaptation (AUFA) framework, which is trained using labeled source domain data and unlabeled target domain data for prediction in the target domain. As shown in Fig. 1, the AUFA comprises four main components: (1) graph representation learning for extracting rs-fMRI features based on Transformer from the constructed brain FCNs, (2) domain adaptation for feature alignment, (3) augmentation-based self-optimization module, and (4) classification.
@@ IV-B1 Graph Representation Learning Module
Graph Construction. The brain is a complex neural network system, wherein different ROIs are responsible for respective functions and interact with each other to coordinate cognitive processes and behaviors. The FCN constructed by analyzing the fMRI time series relationship between different ROIs can effectively capture the complex information of the brain. Each subject’s FCN can be modeled as an undirected graph $G$ = $(V,E)$, where $V$ is the node set, representing different ROIs, and $E$ is the edge set, representing the strength of functional connections between ROI pairs. Since FCN can capture the topology of the brain, it is widely used for computer-aided diagnosis of brain diseases.
In this work, we first construct FCNs for subjects from the source domain and the target domain, respectively. Specifically, we use Pearson correlation (PC) to calculate correlations of rs-fMRI time series, establishing functional connections between ROI pairs. The functional connection $w_{ij}\in\ [-1,1]$ between the $i$-th ROI and the $j$-th ROI is defined as:
@@
Transformer-based Graph Representation Learning. To accurately capture the specific contributions of individual brain regions and learn more discriminative features, we adopted the Transformer encoder for graph representation learning. This strategy utilizes the Transformer’s self-attention mechanism to weigh the importance of different brain regions and obtain the corresponding attention coefficients. Subsequently, based on the attention weights, the feature information of global nodes is aggregated and updated. Thus, it facilitates a deeper understanding of the structure of complex brain networks. More details are introduced in the following.
@@
To capture more comprehensive information about the graph data, we use a multi-head self-attention mechanism for parallel computation. The input $X^{d}$ is passed to $H$ different self-attention modules and $H$ output matrices $Z$ are computed and aggregated. The aggregation operation can be expressed as:
@@
Based on the output of the self-attention layer, a further nonlinear transformation is performed by a feed-forward neural network. The feed-forward neural network is a two-layer fully connected neural network with the following formulation:
${Z^{l+1}}=N(\Psi(\bar{Z}^{l}W_{1}+b_{1})W_{2}+b_{2}),$ (4)
where $\Psi$ denotes the $RELU$ activation function and $W_{1}$, $W_{2}$, $b_{1}$, $b_{2}$ are the learnable weights and biases. Thus, a new feature representation $Z_{X^{d}}$ of $X^{d}$ is obtained through the learning of multiple Transformer layers. To conduct the subsequent graph classification task, the $Z_{X^{d}}$ are flattened to obtain the final graph-level representation $f_{X^{d}}$.

### v2

@@ I Introduction
Figure 1: Overview of our augmentation-based unsupervised cross-domain fMRI adaptation (AUFA) framework, including (1) fMRI feature extraction for brain functional connectivity networks using a Transformer module, (2) domain adaptation module for feature alignment between source and target domains, (3) augmentation-based self-optimization module, and (4) classification module. The data from the target domain are fed into the model that has been jointly optimized by $L_{C}$, $L_{M}$ and $L_{A}$ for direct prediction.
@@
Major depressive disorder (MDD) is a common mental health disorder, leading to symptoms such as profound sadness, self-denial, feelings of helplessness, and impaired cognitive function, thus significantly impacting the patient’s daily life and work [1]. Currently, the diagnosis of MDD mainly relies on the patient’s symptom assessment, physical examination, and psychological scales, which still involves subjectivity and risk of misdiagnosis [2, 3]. Therefore, precise identification of MDD is crucial for clinicians to develop prompt intervention plans.
@@
Resting-state functional magnetic resonance imaging (rs-fMRI) is used as a noninvasive, nonradioactive technique to assess changes in neural activity over time in the whole brain. It is extensively utilized in computer-aided analysis of brain disorders using learning-based methods [4, 5]. Currently, machine learning and deep learning techniques primarily focus on utilizing functional connectivity networks (FCNs) to capture the intricate interplay among regions of interest (ROIs) within the brain [6]. These networks are constructed from blood-oxygen-level dependent (BOLD) time series data obtained from rs-fMRI [7]. Traditional machine learning methods often extract handcrafted features based on FCN for brain disease detection. However, these methods require prior knowledge of FCN [8], and shallow features usually result in poor performance in classification tasks.
@@
With advances in deep learning, various methods employ graph neural networks to automatically extract high-level feature representations from fMRI time series for automatic diagnosis of MDD [9, 10]. For example, Li et al[11] learned topological features from brain functional connectivity constructed based on fMRI data, and utilized a graph neural network (GNN) to identify MDD. Jun et al[12] integrated valid connectivity features from rs-fMRI with non-imaging phenotypic features, followed by training a graph convolutional network (GCN) on the integrated features. However, these methods only use a single site from the MDD dataset for experiments, which does not guarantee the generalization ability of the model. Multi-site rs-fMRI data provide more data for learning-based methods to train reliable models [13, 14]. However, fMRI acquired from different sites often exhibit significant heterogeneity due to equipment differences and variations in data acquisition conditions [15], which presents a challenge for diagnosing brain diseases.
@@
Domain adaptation helps reduce the distributional gap between domains [16]. However, when there is a large discrepancy between the source and target domains, along with the size limitations of fMRI datasets, overfitting on the source domain is more likely to occur [17]. That is, the model is too fine-grained in capturing the source domain features and ignores the generalization of the universal laws. Intuitively, augmenting the target domain data can improve the model’s ability to generalize to the target domain and avoid over-reliance on specific source domain data [18, 19].
@@
To this end, we develop a new augmentation-based unsupervised cross-domain fMRI adaptation (AUFA) framework for automated diagnosis of MDD. As shown in Fig. 1, we first construct FCNs for source/target domain subjects based on fMRI data. Then, we employ the Transformer encoder to learn fMRI features with spatial attention. At the same time, the maximum mean difference constraint is introduced to align the fMRI features learned between domains, aiming for the Transformer encoder to learn shared features between domains. Furthermore, to enhance the model’s generalization ability to the target domain, we optimize it by augmenting target domain data to prevent overfitting on the source domain. The AUFA significantly reduces the data distribution differences across sites through the supervision of source domain labels and adaptation of data between domains. Experiments on rs-fMRI data from 1,089 subjects indicate the excellent performance of AUFA in identifying MDD compared to other competing methods.
The major contributions of this work are summarized as follows.
@@
  • A novel augmentation-based unsupervised cross-domain fMRI adaptation (AUFA) framework is proposed for automated diagnosis with rs-fMRI data.
@@
  • The AUFA introduces the maximum mean difference constraint to align fMRI features from different sites, which can significantly reduce the discrepancy of data distribution between source and target domains.
@@
  • AUFA can effectively capture the attention relationship between different ROIs to automatically learn fMRI features and obtain discriminative functional connectivity.
  • Cross-domain disease detection experiments at 5 fMRI imaging sites demonstrate the superiority of our method over several advanced methods.
The remainder of this paper is structured as follows. A concise overview of relevant studies is provided in Section II. Section III introduces the data acquisition and pre-processing. Section IV introduces the details of the proposed AUFA. We describe the experimental settings, comparison methods, and experimental results in Section V. Section VI discusses the primary components of the proposed AUFA and lists several limitations in the present work. Finally, the paper is summarized in Section VII.
@@ II-A Learning-based Methods for Functional MRI Analysis
Deep learning, as a powerful machine learning method, has been widely used in fMRI analysis [20, 21]. Deep learning models exhibit superior capabilities in handling high-dimensional, nonlinear, and complex fMRI data, showcasing better modeling and generalization performance compared to traditional machine learning methods [22]. Graph neural networks, as a class of deep learning algorithms dedicated to processing graph data, have also been widely used for the analysis of FCN based on fMRI, such as graph convolutional networks (GCNs) [23]. GCNs learn node representations by aggregating features of nodes and their neighbors through convolutional operations. Yao et al[24] developed a mutual multi-scale triplet GCN for diagnosing brain diseases, analyzing both brain functional and structural connectivity. Noman et al[25] designed a new GCN-based graphical autoencoder (GAE) framework for MDD detection. However, these methods usually ignore the extent to which different brain regions contribute to downstream tasks. Transformer is an attention-based deep learning model originally proposed for natural language processing tasks (e.g., machine translation) [26]. It introduces a self-attention mechanism that enables modeling the dependencies between different positions in an input sequence. In recent years, the Transformer model has also been applied as a powerful sequence modeling tool for capturing attention relationships between brain regions in fMRI data analysis [27, 28]. In this work, we attempt to leverage the Transformer in unsupervised domain adaptation for multi-site fMRI feature extraction. Considering that the strong backbone of Transformer increases the risk of overfitting in the source domain, we augment the target domain data and improve the model generalization through a self-optimization module.
@@ II-B Domain Adaptation of Functional MRI Data
The multi-site fMRI data greatly increased the sample size for the model training. However, due to distributional differences among data from different sites, the training data may not precisely represent the distribution of the testing data, leading to a degradation in model performance [29, 15]. The domain adaptation approach is an important strategy for solving this problem. Currently, domain adaptation methods can be categorized into three groups based on the availability of target domain labels [30]: (1) supervised domain adaptation [31], (2) Semi-supervised domain adaptation [32], and (3)unsupervised domain adaptation [33]. In this work, we focus on unsupervised domain adaptation, where only unlabeled target data can be used to train the adaptive model.
Recently, domain adaptation methods have been extensively used in the field of computer vision [34]. For example, Zhou et al[35] proposed a new Adversarial Distribution Adaptation Network (ADAN) to simultaneously minimize global and local distributional differences between domains, facilitating the learning of domain-invariant representations. Li et al[36] proposed a Progressive Adaptation of Subspace (PAS) method to gradually obtain reliable pseudo-labels for model training. Liu et al[37] developed a new unsupervised domain adaptation method that preserves the data distribution of the source domain to guide the learning of target data structure through a semi-supervised clustering approach. However, multi-site adaptation studies for fMRI time series are still limited. Some studies based on multi-site fMRI data have simply combined fMRI data from different sites for training [38], i.e., assuming that they have the same data distribution, ignoring data heterogeneity between sites. Even though some studies attempt to reduce the distributional differences of fMRI data through domain adaptation methods, they usually ignore the problem that the model is likely to be overfitted on the source domain. For example, Wang et al[39] proposed an unsupervised graph domain adaptation network for autism spectrum detection, comprising three components: graph isomorphism encoder, progressive feature alignment, and unsupervised information matrix regularization. Lee et al[29] designed a novel domain generalization meta-learning model to identify brain diseases based on rs-fMRI. The aim of this paper is to study unsupervised domain adaptation for multi-site fMRI data from two perspectives, i.e., reducing inter-domain distributional differences and mitigating the risk of model overfitting.
TABLE I: Phenotypic information of 1089 subjects from REST-meta-MDD consortium [40]. MDD: Major depressive disorder; NC: Normal control; M/F: Male/Female.
@@ III-A Data Acquisition
This work involved rs-fMRI data from the top five largest sites in the imaging center of the REST-meta-MDD consortium [40], including SITE 20, SITE 21, SITE 25, SITE 1, and SITE 9. There were a total of 1,089 participants across the five sites, including 581 MDD patients and 508 normal controls (NCs). The rs-fMRI data from SITE 20, SITE 21 and SITE 1 were acquired by Siemens Verio 3T with the following parameters: echo time (TE)///repetition time (TR) $=30\,ms/2,000\,ms$, flip angle $=90$. Particularly, at SITE 20, the thickness///gap was $3.0\,mm/1.0\,mm$ with $32$ slice; at SITE 21, the thickness///gap was $3.5\,mm/0.7\,mm$ with $32$ slice, and at SITE 1, the thickness///gap was $4.0\,mm/0.8\,mm$ with $32$ slice. The rs-fMRI data from SITE 25 were obtained by Siemens Verio 3T with the following parameters: TE///TR $=25\,ms/2,000\,ms$, flip angle $=90$, thickness///gap $=4.0\,mm/0\,mm$, slice number $=36$. The rs-fMRI data for SITE 9 were obtained by GE Discovery MR750 3.0T with the following parameters: TE///TR $=25\,ms/2,000\,ms$, flip angle $=90$, thickness///gap = $3.0\,mm/1.0\,mm$, slice number $=35$. Table I reports more detailed demographic information.
@@ III-B Image Pre-Processing
The initial rs-fMRI data were pre-processed using the Data Processing Assistant for Resting-State fMRI (DPARSF) [41] software, including the following steps: (1) discard the first ten time points, (2) head motion correction, (3) space normalization and smoothing, (4) interference signal regression, (5) band-pass filtering for noise reduction. Finally, we extracted an average rs-fMRI time series for 116 predefined ROI based on Automated Anatomical Labeling (AAL) [42] atlas.
@@ IV-A Problem Formulation
In this work, we focus on utilizing cross-domain rs-fMRI data for the MDD identification task, aiming to solve an interesting, realistic, but challenging unsupervised domain adaptation problem, i.e., Utilizing knowledge of the source domain to help predict completely unlabeled target domain data. We first introduce some notation that will be used in this study. Let $X^{d}$ be the input (graphs constructed based on fMRI data) space and $Y^{d}$ be the target (label) space, where $d$ identifies either the source domain $s$ or the target domain $t$. The source domain is defined as $S$ = $\{X^{s},Y^{s}\}$, containing the feature set $X^{s}$ and the corresponding label set $Y^{s}$. The target domain is defined accordingly as $T$ = $\{X^{t},Y^{t}\}$ with feature set $X^{t}$ and label set $Y^{t}$. Even though the two domains have the same task (they share the label space $Y^{s}$ = $Y^{t}$), their data distribution is somewhat different, i.e., $P^{s}(X)\neq P^{t}(X)$. In the framework of unsupervised domain adaptation, the challenge is how to learn a feature extractor $\Psi:X\rightarrow C$ that leverages the abundant labeling information from the source domain. This extractor maps the initial input to a feature space $C$, aiming to reduce the data distribution between the source and target domains, i.e., $P^{s}(\Psi(X))\approx P^{t}(\Psi(X))$.
@@ IV-B Method
We propose an augmentation-based unsupervised cross-domain fMRI adaptation (AUFA) framework for MDD identification. As shown in Fig. 1, the AUFA comprises four main components: (1) graph representation learning for extracting rs-fMRI features based on Transformer from the constructed brain FCNs, (2) domain adaptation for feature alignment, (3) augmentation-based self-optimization module, and (4) classification.
@@ IV-B1 Graph Representation Learning Module
Graph Construction. The brain is a complex neural network system, wherein different ROIs are responsible for respective functions and interact with each other to coordinate cognitive processes and behaviors. The FCN constructed by analyzing the fMRI time series relationship between different ROIs can effectively capture the complex relationship in the brain. Each subject’s FCN can be modeled as an undirected graph $G$ = $(V,E)$, where $V$ denotes the node set, representing different ROIs, and $E$ denotes the edge set, representing the strength of functional connections between ROI pairs. FCNs are widely employed in computer-aided detection of MDD due to their ability to capture the topology of the brain.
In this work, we first construct FCNs for subjects from the source and target domains. We utilize Pearson correlation (PC) to calculate correlations of rs-fMRI time series, establishing functional connections between ROI pairs. The functional connection $w_{ij}\in\ [-1,1]$ between the $i$-th ROI and the $j$-th ROI is defined as:
@@
Transformer-based Graph Representation Learning. To accurately capture the specific contributions of individual brain regions and learn more discriminative features, we adopted the Transformer encoder for graph representation learning. This strategy utilizes the Transformer’s self-attention mechanism to weigh the importance of different ROIs and learn the corresponding attention coefficients. Subsequently, based on the attention weights, the feature information of global nodes is aggregated and updated. Thus, it facilitates a deeper understanding of the structure of complex brain networks. More details are introduced in the following.
@@
To capture more comprehensive information about the graph data, we employ a multi-head self-attention mechanism for parallel computation. The input $X^{d}$ is passed to $H$ different self-attention modules and $H$ output matrices $Z$ are computed and aggregated. The aggregation operation can be represented as:
@@
After feature learning through self-attention layers , a further nonlinear transformation is performed by a feed-forward neural network. The feed-forward neural network is a two-layer fully connected neural network with the following formulation:
${Z^{l+1}}=N(RELU(\bar{Z}^{l}W_{1}+b_{1})W_{2}+b_{2}),$ (4)
where $W_{1}$, $W_{2}$, $b_{1}$, $b_{2}$ are the learnable weights and biases. Thus, a new feature representation $Z_{X^{d}}$ of $X^{d}$ is obtained through the learning of multiple Transformer layers. To conduct the subsequent graph classification task, the $Z_{X^{d}}$ are flattened to obtain the final graph-level representation $f_{X^{d}}$.